Mercer Asbestos Removal Company removes potentially toxic asbestos insulation and related products from buildings. There has been a long-simmering dispute between the company’s estimator and the work supervisors. The on-site supervisors claim that the estimators do not adequately distinguish between routine work, such as removal of asbestos insulation around heating pipes in older homes, and nonroutine work, such as removing asbestos-contaminated ceiling plaster in industrial buildings. The on-site supervisors believe that nonroutine work is far more expensive than routine work and should bear higher customer charges. The estimator sums up his position in this way: “My job is to measure the area to be cleared of asbestos. As directed by top management, I simply multiply the square footage by $3.20 to determine the bid price. Since our average cost is only $2.745 per square foot, that leaves enough cushion to take care of the additional costs of nonroutine work that shows up. Besides, it is difficult to know what is routine or not routine until you actually start tearing things apart.”

To shed light on this controversy, the company initiated an activity-based costing study of all of its costs. Data from the activity-based costing system follow:

Activity Cost Pool Activity Measure Total Activity
Removing asbestos Thousands of square feet 800 thousand square feet
Estimating and job setup Number of jobs 500 jobs
Working on nonroutine jobs Number of nonroutine jobs 100 nonroutine jobs
Other (organization-sustaining costs and idle capacity costs) None
Note: The 100 nonroutine jobs are included in the total of 500 jobs. Both nonroutine jobs and routine jobs require estimating and setup.

Costs for the Year

Wages and salaries $ 470,000
Disposal fees 801,000
Equipment depreciation 110,000
On-site supplies 68,000
Office expenses 370,000
Licensing and insurance 570,000
Total cost $ 2,389,000
Distribution of Resource Consumption Across Activities
Removing Asbestos Estimating and Job Setup Working on Nonroutine Jobs Other Total
Wages and salaries 60 % 10 % 20 % 10 % 100 %
Disposal fees 60 % 0 % 40 % 0 % 100 %
Equipment depreciation 50 % 10 % 15 % 25 % 100 %
On-site supplies 60 % 25 % 15 % 0 % 100 %
Office expenses 10 % 40 % 15 % 35 % 100 %
Licensing and insurance 30 % 0 % 50 % 20 % 100 %

1. Perform the first-stage allocation of costs to the activity cost pools.
Answer #1


Mercer Asbestos Removal Company Distribution of Resource Consumption Across Activities Other 47000 Removing Estimating & Asbe

1 Mercer Asbestos Removal Cor 2 3 Distribution of Resource Cons 5 Estimating & Job Setup =F7*0.1 Other =F7*0.1 0 Removing Asb